# LLM Unlearning with LLM Beliefs

### v1

@@ 6 Conclusions
Our research directly addresses ethical concerns related to harmful knowledge in LLMs. We propose methods to reliably remove undesirable information, reducing risks of privacy violations and harmful content exposure. Experiments utilized public datasets without direct human involvement, mitigating privacy risks. Methodological limitations and potential risks are transparently reported to promote trust and ongoing improvement in AI systems.
@@
We ensure reproducibility by clearly documenting experimental setups, methods, benchmarks, model architectures, and hyperparameters for proposed methods. Complete theoretical proofs are provided in the appendix. Source code will be released after published.
@@
  • Appendix A introduces the notations used throughout the paper.
  • Appendix B expands the related work discussion, covering both machine unlearning in general and LLM unlearning in particular.
  • Appendix C presents pseudocode for BS-T and BS-S, clarifying their implementation.
  • Appendix D elaborates the theoretical framework, including the learning dynamics perspective and the full proof of Theorem 4.2.
  • Appendix E details the experimental setup, including benchmarks, evaluation metrics, implementation configurations, and hyperparameters.
  • Appendix F reports additional results: Appendix F.1 provides more failure case examples; Appendix F.2 introduces the prompt design for LaaJ evaluation; Appendix F.3 gives results on MUSE-Books; Appendix F.4 shows qualitative examples of unlearning responses; Appendix F.5 conducts ablation studies; and Appendix F.6 summarizes time consumption.
  • Appendix G discusses the limitations of our approach.
  • Appendix H outlines the usage of LLMs in this work.
@@ Appendix A Notations
In this section, we summarize important notations in Table 4.
@@ Appendix B Detailed Related Works
This section reviews the literature of machine unlearning in Appendix B.1 and LLM unlearning in Appendix B.2.
@@ B.1 Machine Unlearning
Early studies on machine unlearning aim to enable the removal of specific training data post hoc, motivated by privacy regulations like the “right to be forgotten” (Xu et al., 2024). Cao & Yang (2015) first introduce the concept by reformulating algorithms (e.g., naïve Bayes, k-means, SVM) to support exact unlearning via efficient updates without full retraining. Ginart et al. (2019) later formalize data deletion as updating a trained model to behave as if the data were never seen, proposing efficient deletion algorithms to avoid redundant computations. Subsequent work distinguishes between certified and approximate unlearning (Thudi et al., 2022).
Certified methods provide formal guarantees. Guo et al. (2020) leverage influence functions and differential privacy for provable deletion in convex models. Bourtoule et al. (2021) propose SISA (Sharded, Isolated, Sliced, Aggregated training), enabling certified removal by retraining only affected model shards, significantly improving efficiency. Approximate methods trade guarantees for scalability. Influence-based techniques estimate and subtract a data point’s effect (Koh & Liang, 2017; Guo et al., 2020), extended to deep models via approximations (Sekhari et al., 2021; Suriyakumar & Wilson, 2022; Mehta et al., 2022). Though lacking indistinguishability guarantees, these approaches enable fast and practical unlearning. Other approximate methods include noise injection and distillation (Golatkar et al., 2020), domain-specific strategies for forests (Brophy & Lowd, 2021), GNNs (Chen et al., 2022b), and recommender systems (Chen et al., 2022a). Recent efforts extend unlearning to generative models, where knowledge is more entangled. Wu et al. (2020) introduce DeltaGrad for efficient retraining, applicable to generative tasks. Golatkar et al. (2020) explore unlearning in vision models via activation scrubbing. For diffusion models, Chen et al. (2025) propose Score Forgetting Distillation (SFD), aligning generative scores of “forbidden” and “safe” concepts to erase targeted content. Li et al. (2024a) further develop the first method for image-to-image unlearning, enabling removal of specific patterns (e.g., copyright or violence) within seconds.
@@ B.2 LLM Unlearning
One line of work aims to refine unlearning objectives or loss weighting to mitigate collateral damage. Negative Preference Optimization (NPO) reframes unlearning as inverse preference tuning: Zhang et al. (2024) penalize high-probability “undesirable” responses by reweighting gradients, inspired by DPO (Rafailov et al., 2024) but with reversed rewards. Fan et al. (2025) simplify NPO (dubbed SimNPO) by smoothing its gradient weighting scheme, and report more stable forgetting behavior. Mekala et al. (2025) propose an alternate preference optimization (AltPO) mechanism as another variant to improve stability during unlearning. Beyond preference-based objectives, several works adjust loss contributions dynamically. Wang et al. (2025b) propose weighted gradient ascent (WGA) that leverages the conditional token form of GA, further incorporating token-wise weighting to enable more fine-grained control. Yang et al. (2025) advocate focusing on under-forgotten examples (“saturation” perspective) versus high-impact examples (“importance”), designing a an improved loss reweighting method named Saturation-Importance (SatImp) to optimize the forget–retain trade-off. Wang et al. (2025c) even show that tuning only on the forget set—without explicit retain data—can still preserve utility with proper scheduling. Other techniques include label smoothing for forget-set targets and directly optimizing logit differences between forget and retain outputs. To mitigate side effects, Wang et al. (2025d) introduce Gradient Rectified Unlearning (GRU) to explicitly preserve performance during forgetting. Altogether, these methods improve over naïve GA by modulating training signals—via weighting, loss shaping, or multi-objective design—for a more balanced forget–retain trade-off.
@@
Another line of work alters which model parameters or representations are updated during unlearning. Rather than tuning all weights, parameter-efficient unlearning methods add small modulatory components to the model. Chen & Yang (2023) insert “unlearning layers” (lightweight adapter modules) into the transformer; these are trained with a selective teacher–student objective to forget specified data without disturbing other knowledge. This approach allows quick updates and even supports sequential unlearning: different unlearning layers can be fused to handle multiple forget requests. Bhaila et al. (2025) take a prompting approach, learning soft prompts that, when prepended, suppress specific knowledge without altering the base model. Liu et al. (2024) propose embedding-corrupted prompts that inject adversarial noise into token embeddings, disrupting internal representations of the forget set. Other methods intervene at the representation level: Li et al. (2024b) introduce Representation Misdirection Unlearning (RMU), which perturbs hidden activations linked to the forget set to erase associated outputs. Shen et al. (2025) propose LUNAR, which operates unlearning by redirecting the representations of unlearned data to regions that trigger the model’s inherent ability to express its inability to answer.
@@ 14    ${\bm{\theta}}\leftarrow{\bm{\theta}}-\eta\,\nabla_{\bm{\theta}}\mathcal{L}$;
Appendix D Theoretical Framework and Missing Proofs
This section formalizes how the learning–dynamics framework of Ren & Sutherland (2025) can be re-used to analyze unlearning objectives. We first recap the decomposition for one-step prediction change in LLM finetuning (Appendix D.1), then instantiate it for GA and BS-T to prove Theorem 4.2 and discuss its implications (Appendix D.2).
@@ D.1 Learning Dynamics
We work at a single step $t\to t+1$ of SGD on one unlearning sample $\bm{\mathsf{\chi}}_{\rm u}=[{\mathbf{x}}_{\rm u};{\mathbf{y}}_{\rm u}]$ with learning rate $\eta$. Let $h_{\theta}(\bm{\mathsf{\chi}})={\mathbf{z}}(\bm{\mathsf{\chi}})\in\mathbb{R}^{V\times L}$ be the token–logit matrix (sequence length $L$) and $\pi_{\bm{\theta}}(\cdot|\bm{\mathsf{\chi}})=\mathrm{softmax}({\mathbf{z}}(\bm{\mathsf{\chi}}))$ the conditional distribution under teacher forcing, which lets us index tokens by $\bm{\mathsf{\chi}}=[{\mathbf{x}};{\mathbf{y}}]$ and treat the autoregressive structure through the causal mask of $h_{\bm{\theta}}$. [^2] Teacher forcing makes ${\mathbf{y}}^{<i}$ given rather than sampled, which allows bundling an input–output pair into $\bm{\mathsf{\chi}}$ and using a shared kernel $\mathcal{K}_{t}(\bm{\mathsf{\chi}}_{\rm o},\bm{\mathsf{\chi}}_{\rm u})$ for sequence models.
@@
Define the empirical NTK (for the logit network) $\mathcal{K}_{t}(\bm{\mathsf{\chi}}_{\rm o},\bm{\mathsf{\chi}}_{\rm u})=\nabla_{\bm{\theta}}{\mathbf{z}}(\bm{\mathsf{\chi}}_{\rm o})\nabla_{\bm{\theta}}^{\top}{\mathbf{z}}(\bm{\mathsf{\chi}}_{\rm u})$ and the residual $\mathcal{G}_{t}(\bm{\mathsf{\chi}}_{\rm u})=\nabla_{{\mathbf{z}}}\mathcal{L}(\bm{\mathsf{\chi}}_{\rm u})$; the $\mathcal{O}(\eta^{2})$ remainder follows the same bound as in Ren & Sutherland (2025).
@@
and let $t^{i}\in\Delta^{|\mathcal{V}|-1}$ be a generic target distribution. The token-wise negative cross-entropy is
$\mathcal{L}(\bm{\mathsf{\chi}}_{\rm u})\;=\;-\sum_{i=1}^{L}\big\langle t^{i},\log\pi^{i}\big\rangle\;=\;-\sum_{i=1}^{L}\sum_{v\in\mathcal{V}}t^{i}[v]\log\pi^{i}[v].$ (8)
@@
$\frac{\partial\,\log\pi^{i}[v]}{\partial\,{\mathbf{z}}^{i}[r]}\;=\;\mathbf{1}\{v=r\}\;-\;\frac{e^{{\mathbf{z}}^{i}[r]}}{\sum_{u}e^{{\mathbf{z}}^{i}[u]}}\;=\;\mathbf{1}\{v=r\}-\pi^{i}[r].$ (9)
@@
$\frac{\partial\,\mathcal{L}(\bm{\mathsf{\chi}}_{\rm u})}{\partial\,{\mathbf{z}}^{i}[r]}\;=\;-\sum_{v\in\mathcal{V}}t^{i}[v]\,\frac{\partial\,\log\pi^{i}[v]}{\partial\,{\mathbf{z}}^{i}[r]}\;=\;-\sum_{v}t^{i}[v]\big(\mathbf{1}\{v=r\}-\pi^{i}[r]\big)\;=\;-t^{i}[r]+\Big(\sum_{v}t^{i}[v]\Big)\pi^{i}[r].$
Since $t^{i}\in\Delta^{|\mathcal{V}|-1}$, $\sum_{v}t^{i}[v]=1$, hence
$\nabla_{{\mathbf{z}}^{i}}\mathcal{L}(\bm{\mathsf{\chi}}_{\rm u})\;=\;\pi^{i}-t^{i}.$ (10)
For GA, $t^{i}={\mathbf{e}}_{y^{i}_{\rm u}}$, so by Eq. (10)
@@
so $\partial{\mathbf{q}}^{i}/\partial{\mathbf{z}}^{i}=\mathbf{0}$. BS-T uses the convex target
$t^{i}_{\rm BST}\;=\;(1-\lambda)\,{\mathbf{e}}_{y^{i}_{\rm u}}\;+\;\lambda\,{\mathbf{q}}^{i},$
whence, by (10) and the stop-gradient property of ${\mathbf{q}}^{i}$,
$\mathcal{G}^{i}_{\rm BST}\;=\;\nabla_{{\mathbf{z}}^{i}}\mathcal{L}(\bm{\mathsf{\chi}}_{\rm u})\;=\;\pi^{i}-t^{i}_{\rm BST}\;=\;\pi^{i}-\Big((1-\lambda)\,{\mathbf{e}}_{y^{i}_{\rm u}}+\lambda\,{\mathbf{q}}^{i}\Big).$

### v2

@@ 7 Conclusions
Acknowledgment
This work was supported in part by Macau Science and Technology Development Fund under 001/2024/SKL, 0119/2024/RIB2, 0110/2025/R1B2, and 0022/2022/A1; in part by Research Committee at University of Macau under MYRG-CRG2025-00031-FST and MYRG-GRG2025-00086-FST; in part by the Guangdong Basic and Applied Basic Research Foundation under Grant 2024A1515012536; in part by RGC General Research Fund No. 12200725 and RGC Young Collaborative Research Grant No. C2005-24Y.
@@ Acknowledgment
In accordance with the ICLR Code of Ethics, our research directly addresses ethical concerns related to harmful knowledge in LLMs. We propose methods to reliably remove undesirable information, reducing risks of privacy violations and harmful content exposure. Experiments utilized public datasets without direct human involvement, mitigating privacy risks. Methodological limitations and potential risks are transparently reported to promote trust and ongoing improvement in AI systems.
@@
We ensure reproducibility by clearly documenting experimental setups, methods, benchmarks, model architectures, and hyperparameters for proposed methods. Complete theoretical proofs are provided in the appendix, with code merged to OpenUnlearning.
Usage of Large Language Models
@@
  • §A introduces the notations used throughout the paper.
  • §B expands the related work discussion, covering both machine unlearning in general and LLM unlearning in particular.
  • §C presents pseudocode for BS-T and BS-S, clarifying their implementation.
  • §D provides the theoretical details, including the AKG decomposition and the proofs of Thms. 5.2 and 5.3, together with a brief discussion of the on-policy case.
  • §E details the experimental setup, including benchmarks, evaluation metrics, implementation configurations, and hyperparameters.
  • §F reports additional results: §F.1 provides more failure case examples; §F.2 introduces the prompt design for LaaJ evaluation; §F.3 gives results on MUSE; §F.4 shows qualitative examples of unlearning responses; §F.5 conducts ablation studies; and §F.6 summarizes time consumption.
  • §G discusses the limitations of our approach.
@@ Appendix A Notations
In this section, we summarize important notations in Tab. 3.
@@ Appendix B Detailed Related Works
This section reviews the literature of machine unlearning in §B.1 and LLM unlearning in §B.2.
@@ B.1 Machine Unlearning
Early studies on machine unlearning aim to enable the removal of specific training data post hoc, motivated by privacy regulations like the “right to be forgotten” (Xu et al., 2024). Cao and Yang (2015) first introduce the concept by reformulating algorithms (e.g., naïve Bayes, k-means, SVM) to support exact unlearning via efficient updates without full retraining. Ginart et al. (2019) later formalize data deletion as updating a trained model to behave as if the data were never seen, proposing efficient deletion algorithms to avoid redundant computations. Subsequent work distinguishes between certified and approximate unlearning (Thudi et al., 2022).
Certified methods provide formal guarantees. Guo et al. (2020) leverage influence functions and differential privacy for provable deletion in convex models. Bourtoule et al. (2021) propose SISA (Sharded, Isolated, Sliced, Aggregated training), enabling certified removal by retraining only affected model shards, significantly improving efficiency. Approximate methods trade guarantees for scalability. Influence-based techniques estimate and subtract a data point’s effect (Koh and Liang, 2017; Guo et al., 2020), extended to deep models via approximations (Sekhari et al., 2021; Suriyakumar and Wilson, 2022; Mehta et al., 2022). Though lacking indistinguishability guarantees, these approaches enable fast and practical unlearning. Other approximate methods include noise injection and distillation (Golatkar et al., 2020), domain-specific strategies for forests (Brophy and Lowd, 2021), GNNs (Chen et al., 2022b), and recommender systems (Chen et al., 2022a). Recent efforts extend unlearning to generative models, where knowledge is more entangled (Li et al., 2026). Wu et al. (2020) introduce DeltaGrad for efficient retraining, applicable to generative tasks. Golatkar et al. (2020) explore unlearning in vision models via activation scrubbing. For diffusion models, Chen et al. (2025) propose Score Forgetting Distillation (SFD), aligning generative scores of “forbidden” and “safe” concepts to erase targeted content. Li et al. (2024a) further develop the first method for image-to-image unlearning, enabling removal of specific patterns (e.g., copyright or violence) within seconds. Other works such as Yu et al. (2025) customize unlearning method for multi-task learning, and Luo et al. (2026) apply machine unlearning to GNN for social networks.
@@ B.2 LLM Unlearning
One line of work aims to refine unlearning objectives or loss weighting to mitigate collateral damage. Negative Preference Optimization (NPO) reframes unlearning as inverse preference tuning: Zhang et al. (2024) penalize high-probability “undesirable” responses by reweighting gradients, inspired by DPO (Rafailov et al., 2024) but with reversed rewards. Fan et al. (2025) simplify NPO (dubbed SimNPO) by smoothing its gradient weighting scheme, and report more stable forgetting behavior. Mekala et al. (2025) propose an alternate preference optimization (AltPO) mechanism as another variant to improve stability during unlearning. Beyond preference-based objectives, several works adjust loss contributions dynamically. Wang et al. (2025b) propose weighted gradient ascent (WGA) that leverages the conditional token form of GA, further incorporating token-wise weighting to enable more fine-grained control. Yang et al. (2025) advocate focusing on under-forgotten examples (“saturation” perspective) versus high-impact examples (“importance”), designing a an improved loss reweighting method named Saturation–Importance (SatImp) to optimize the forget–retain trade-off. Wang et al. (2025c) even show that tuning only on the forget set—without explicit retain data—can still preserve utility with proper scheduling. Other techniques include label smoothing for forget-set targets and directly optimizing logit differences between forget and retain outputs. To mitigate side effects, Wang et al. (2025d) introduce Gradient Rectified Unlearning (GRU) to explicitly preserve performance during forgetting. Altogether, these methods improve over naïve GA by modulating training signals—via weighting, loss shaping, or multi-objective design—for a more balanced forget–retain trade-off.
@@
Another line of work alters which model parameters or representations are updated during unlearning. Rather than tuning all weights, parameter-efficient unlearning methods add small modulatory components to the model. Chen and Yang (2023) insert “unlearning layers” (lightweight adapter modules) into the transformer; these are trained with a selective teacher–student objective to forget specified data without disturbing other knowledge. This approach allows quick updates and even supports sequential unlearning: different unlearning layers can be fused to handle multiple forget requests. Bhaila et al. (2025) take a prompting approach, learning soft prompts that, when prepended, suppress specific knowledge without altering the base model. Liu et al. (2024) propose embedding-corrupted prompts that inject adversarial noise into token embeddings, disrupting internal representations of the forget set. Other methods intervene at the representation level: Li et al. (2024b) introduce Representation Misdirection Unlearning (RMU), which perturbs hidden activations linked to the forget set to erase associated outputs. Shen et al. (2025) propose LUNAR, which operates unlearning by redirecting the representations of unlearned data to regions that trigger the model’s inherent ability to express its inability to answer.
@@ 14    ${\bm{\theta}}\leftarrow{\bm{\theta}}-\eta\,\nabla_{\bm{\theta}}\mathcal{L}$;
Appendix D Theoretical Proofs and Discussions
This section contains the theoretical components supporting §5. §D.1 presents the AKG one-step decomposition of Ren and Sutherland (2025) and proves Lem. 5.1. §D.2 applies this decomposition to GA and BS-T and proves Thm. 5.2. §D.3 extends the analysis to off-policy BS-S and proves Thm. 5.3. §D.4 discusses why the on-policy BS-S variant falls outside this framework.
@@ D.1 Learning Dynamics
We work at a single step $t\to t+1$ of SGD on one unlearning sample $\bm{\mathsf{\chi}}_{\rm u}=[{\mathbf{x}}_{\rm u};{\mathbf{y}}_{\rm u}]$ with learning rate $\eta$. Let $h_{\bm{\theta}}(\bm{\mathsf{\chi}})={\mathbf{z}}(\bm{\mathsf{\chi}})\in\mathbb{R}^{V\times L}$ be the token–logit matrix (sequence length $L$) and $\pi_{\bm{\theta}}(\cdot|\bm{\mathsf{\chi}})=\mathrm{softmax}({\mathbf{z}}(\bm{\mathsf{\chi}}))$ the conditional distribution under teacher forcing, which lets us index tokens by $\bm{\mathsf{\chi}}=[{\mathbf{x}};{\mathbf{y}}]$ and treat the autoregressive structure through the causal mask of $h_{\bm{\theta}}$. [^2] Teacher forcing makes ${\mathbf{y}}^{<i}$ given rather than sampled, which allows bundling an input–output pair into $\bm{\mathsf{\chi}}$ and using a shared kernel $\mathcal{K}_{t}(\bm{\mathsf{\chi}}_{\rm o},\bm{\mathsf{\chi}}_{\rm u})$ for sequence models.
@@
Define the empirical NTK (for the logit network) $\mathcal{K}_{t}(\bm{\mathsf{\chi}}_{\rm o},\bm{\mathsf{\chi}}_{\rm u})=\nabla_{\bm{\theta}}{\mathbf{z}}(\bm{\mathsf{\chi}}_{\rm o})\nabla_{\bm{\theta}}^{\top}{\mathbf{z}}(\bm{\mathsf{\chi}}_{\rm u})$ and the residual $\mathcal{G}_{t}(\bm{\mathsf{\chi}}_{\rm u})=\nabla_{{\mathbf{z}}}\mathcal{L}(\bm{\mathsf{\chi}}_{\rm u})$; the $\mathcal{O}(\eta^{2})$ remainder follows the same bound as in Ren and Sutherland (2025).
@@
and let ${\mathbf{t}}^{i}\in\Delta^{|\mathcal{V}|-1}$ be a generic target distribution. The token-wise negative cross-entropy is
$\mathcal{L}(\bm{\mathsf{\chi}}_{\rm u})\;=\;-\sum_{i=1}^{L}\big\langle{\mathbf{t}}^{i},\log\pi^{i}\big\rangle\;=\;-\sum_{i=1}^{L}\sum_{v\in\mathcal{V}}{\mathbf{t}}^{i}[v]\log\pi^{i}[v].$ (8)
@@
$\frac{\partial\,\log\pi^{i}[v]}{\partial\,{\mathbf{z}}^{i}[r]}\;=\;\mathbbm{1}\{v=r\}\;-\;\frac{e^{{\mathbf{z}}^{i}[r]}}{\sum_{u}e^{{\mathbf{z}}^{i}[u]}}\;=\;\mathbbm{1}\{v=r\}-\pi^{i}[r].$ (9)
@@
$\frac{\partial\,\mathcal{L}(\bm{\mathsf{\chi}}_{\rm u})}{\partial\,{\mathbf{z}}^{i}[r]}\;=\;-\sum_{v\in\mathcal{V}}{\mathbf{t}}^{i}[v]\,\frac{\partial\,\log\pi^{i}[v]}{\partial\,{\mathbf{z}}^{i}[r]}\;=\;-\sum_{v}{\mathbf{t}}^{i}[v]\big(\mathbbm{1}\{v=r\}-\pi^{i}[r]\big)\;=\;-{\mathbf{t}}^{i}[r]+\Big(\sum_{v}{\mathbf{t}}^{i}[v]\Big)\pi^{i}[r].$
Since ${\mathbf{t}}^{i}\in\Delta^{|\mathcal{V}|-1}$, $\sum_{v}{\mathbf{t}}^{i}[v]=1$, hence
$\nabla_{{\mathbf{z}}^{i}}\mathcal{L}(\bm{\mathsf{\chi}}_{\rm u})\;=\;\pi^{i}-{\mathbf{t}}^{i}.$ (10)
For GA, ${\mathbf{t}}^{i}={\mathbf{e}}_{y^{i}_{\rm u}}$, so by Eq. (10),
@@
so $\nabla_{{\mathbf{z}}^{i}}{\mathbf{q}}^{i}=\mathbf{0}$. BS-T uses the convex target
${\mathbf{t}}^{i}_{\rm BST}\;=\;(1-\lambda)\,{\mathbf{e}}_{y^{i}_{\rm u}}\;+\;\lambda\,{\mathbf{q}}^{i},$
whence, by Eq. (10) and the stop-gradient property of ${\mathbf{q}}^{i}$,
$\mathcal{G}^{i}_{\rm BST}\;=\;\nabla_{{\mathbf{z}}^{i}}\mathcal{L}(\bm{\mathsf{\chi}}_{\rm u})\;=\;\pi^{i}-{\mathbf{t}}^{i}_{\rm BST}\;=\;\pi^{i}-\Big((1-\lambda)\,{\mathbf{e}}_{y^{i}_{\rm u}}+\lambda\,{\mathbf{q}}^{i}\Big).$